# JuDGE: Benchmarking Judgment Document Generation for Chinese Legal System

### v1

@@ 5.1.3. Retrieval Augmented Generation
QW-3B-Base SFT 0.5975 0.5149 0.9381 0.9375 0.9378 0.6915 0.6441 0.6669 0.6134 0.7099 0.8068 0.7569
RAG 0.6273 0.5132 0.9471 0.9511 0.9491 0.7569 0.6847 0.7190 0.5811 0.7921 0.8041 0.8512
QW-7B-Base SFT 0.6380 0.5273 0.9651 0.9664 0.9657 0.7500 0.7187 0.7340 0.6413 0.7620 0.8346 0.8163
RAG 0.6489 0.5458 0.9411 0.9441 0.9426 0.7356 0.7649 0.7502 0.5869 0.8206 0.8036 0.8774
LexiLaw-6B Direct 0.0011 0.0000 0.5379 0.5439 0.5409 0.0013 0.0040 0.0020 0.0000 0.0000 0.0000 0.0000
ICL 0.0100 0.0126 0.4072 0.4112 0.4092 0.0070 0.0096 0.0081 0.1846 0.0046 0.1491 0.0057
SFT 0.5855 0.4942 0.9391 0.9295 0.9343 0.6074 0.7018 0.6512 0.5611 0.6564 0.6428 0.6763
Hanfei-7B Direct 0.0691 0.0661 0.1277 0.1281 0.1279 0.0875 0.1053 0.0956 0.4265 0.0781 0.4731 0.0856
ICL 0.0628 0.0282 0.1128 0.1118 0.1123 0.0221 0.0440 0.0295 0.3341 0.0275 0.3568 0.0437
SFT 0.6464 0.5526 0.9381 0.9318 0.9350 0.7298 0.6905 0.7096 0.6172 0.7491 0.8043 0.8045
QW-3B-Chat Direct 0.6117 0.4868 0.8713 0.8752 0.8732 0.5972 0.7300 0.6570 0.4975 0.5815 0.5490 0.7345
ICL 0.6278 0.4720 0.8832 0.8746 0.8789 0.4928 0.7092 0.5815 0.4132 0.6084 0.5983 0.7993
SFT 0.6673 0.5443 0.9401 0.9444 0.9423 0.7205 0.7503 0.7351 0.5717 0.8250 0.7912 0.8923
RAG 0.6527 0.5296 0.9511 0.9531 0.9521 0.7292 0.7484 0.7387 0.5654 0.8345 0.7991 0.8980
QW-7B-Chat Direct 0.6655 0.5075 0.9242 0.9242 0.9242 0.5045 0.8161 0.6235 0.4686 0.7202 0.5537 0.8062
ICL 0.6731 0.5095 0.9371 0.9385 0.9378 0.6537 0.7915 0.7161 0.5406 0.8030 0.6925 0.8534
SFT 0.6604 0.5506 0.9521 0.9518 0.9519 0.7787 0.7694 0.7740 0.6096 0.8421 0.8143 0.9089
RAG 0.6795 0.5512 0.9611 0.9604 0.9607 0.7955 0.7509 0.7726 0.5926 0.8395 0.8116 0.9066
@@ 6. Experimental Results
(1) Legal LLMs (e.g., LexiLaw-6B and Hanfei-7B) show notably poor performance when generating judgment documents directly or via few-shot in-context learning. Even with in-context learning, the performance remains weak. A closer examination of the generated outputs reveals that these models often reuse or partially copy the exemplars provided in the prompt, neglecting to incorporate the unique facts of each case. We suspect that this phenomenon arises from catastrophic forgetting. Since these models were continually pre-trained and fine-tuned on large-scale legal data (e.g., legal QA), their parameters have diverged significantly from their initial state, resulting in the loss of previously acquired knowledge and capabilities. (2) Few-shot in-context learning benefits larger chat models more significantly. For QW-7B-Chat, introducing a few-shot example in the prompt consistently enhances performance across almost all the metrics. In contrast, QW-3B-Chat sees only marginal gains from in-context learning. This difference highlights that larger models with more capacity generally exhibit stronger in-context learning capabilities. (3) Supervised fine-tuning substantially improves performance. Regardless of the models’ initial capability, supervised fine-tuning (SFT) leads to improvements over direct generation and in-context learning. This trend is observed across both general-purpose models (e.g., QW-3B-Chat and QW-7B-Chat) and legal LLMs (e.g., LexiLaw-6B and Hanfei-7B). Notably, the improvement is particularly large for legal LLMs: while their performance is almost unusable before fine-tuning, it becomes acceptable and sometimes competitive after the task-specific SFT. These findings reinforce the importance of supervised adaptation for certain downstream tasks. (4) Multi-source retrieval-augmented generation (RAG) offers competitive advantages. For instance, MRAG consistently achieves the highest recall, precision, and F1 score in convicting accuracy. However, while MRAG achieves high recall in referencing accuracy, its precision is relatively lower. We attribute this to our top-$k$ retrieval strategy (here $k=10$), which occasionally supplies multiple irrelevant statutes. As indicated by Table 2, only 4.31 criminal statutes are truly relevant per case. Since current LLMs are not fully adept at filtering out misleading articles, this results in a precision drop.

### v2

@@ 5.1.3. Retrieval Augmented Generation
QW-3B-Base SFT 0.5975 0.5149 0.9381 0.9375 0.9378 0.6915 0.6441 0.6669 0.6281 0.8400 0.7147 0.7408
RAG 0.6273 0.5132 0.9471 0.9511 0.9491 0.7569 0.6847 0.7190 0.5945 0.8412 0.7045 0.8307
QW-7B-Base SFT 0.6380 0.5273 0.9651 0.9664 0.9657 0.7500 0.7187 0.7340 0.6523 0.8527 0.7548 0.8025
RAG 0.6489 0.5458 0.9411 0.9441 0.9426 0.7356 0.7649 0.7502 0.6020 0.8533 0.7047 0.8565
LexiLaw-6B Direct 0.0408 0.0261 0.6178 0.6228 0.6202 0.0010 0.0040 0.0016 0.0994 0.7259 0.3191 0.0043
ICL 0.0102 0.0126 0.4072 0.4112 0.4092 0.0071 0.0088 0.0078 0.3403 0.7179 0.1636 0.0058
SFT 0.5926 0.4942 0.9401 0.9288 0.9344 0.6030 0.6908 0.6439 0.6471 0.8399 0.6907 0.6715
Hanfei-7B Direct 0.4902 0.3541 0.8623 0.8683 0.8653 0.4624 0.5306 0.4941 0.5132 0.7643 0.5433 0.4860
ICL 0.0635 0.0282 0.1128 0.1118 0.1123 0.0209 0.0407 0.0276 0.3184 0.5404 0.2632 0.0434
SFT 0.6520 0.5526 0.9381 0.9318 0.9350 0.7238 0.6813 0.7019 0.6312 0.8389 0.7141 0.7865
QW-3B-Chat Direct 0.6117 0.4868 0.8713 0.8752 0.8732 0.5972 0.7300 0.6570 0.4620 0.7724 0.3658 0.7120
ICL 0.6278 0.4720 0.8832 0.8746 0.8789 0.4928 0.7092 0.5815 0.4245 0.6901 0.3699 0.6790
SFT 0.6673 0.5443 0.9401 0.9444 0.9423 0.7205 0.7503 0.7351 0.5836 0.8509 0.6803 0.8889
RAG 0.6527 0.5296 0.9511 0.9531 0.9521 0.7292 0.7484 0.7387 0.5807 0.8425 0.7095 0.8851
QW-7B-Chat Direct 0.6655 0.5075 0.9242 0.9242 0.9242 0.5045 0.8161 0.6235 0.5061 0.8098 0.4161 0.7823
ICL 0.6731 0.5095 0.9371 0.9385 0.9378 0.6537 0.7915 0.7161 0.5436 0.8245 0.6220 0.8523
SFT 0.6604 0.5506 0.9521 0.9518 0.9519 0.7787 0.7694 0.7740 0.6220 0.8597 0.7242 0.8961
RAG 0.6795 0.5512 0.9611 0.9604 0.9607 0.7955 0.7509 0.7726 0.6076 0.8513 0.7236 0.8887
@@ 6. Experimental Results
(1) Legal LLMs (e.g., LexiLaw-6B and Hanfei-7B) show notably poor performance when generating judgment documents directly or via few-shot in-context learning. Even with in-context learning, the performance remains weak. A closer examination of the generated outputs reveals that these models often reuse or partially copy the exemplars provided in the prompt, neglecting to incorporate the unique facts of each case. We suspect that this phenomenon arises from catastrophic forgetting. Since these models were continually pre-trained and fine-tuned on large-scale legal data (e.g., legal QA), their parameters have diverged significantly from their initial state, resulting in the loss of previously acquired knowledge and instruction-following capabilities. (2) Few-shot in-context learning benefits larger chat models more significantly. For QW-7B-Chat, introducing a few-shot example in the prompt consistently enhances performance across almost all the metrics. In contrast, QW-3B-Chat sees only marginal gains from in-context learning. This difference highlights that larger models with more capacity generally exhibit stronger in-context learning capabilities. (3) Supervised fine-tuning substantially improves performance. Regardless of the models’ initial capability, supervised fine-tuning (SFT) leads to improvements over direct generation and in-context learning. This trend is observed across both general-purpose models (e.g., QW-3B-Chat and QW-7B-Chat) and legal LLMs (e.g., LexiLaw-6B and Hanfei-7B). Notably, the improvement is particularly large for legal LLMs: while their performance is almost unusable before fine-tuning, it becomes acceptable and sometimes competitive after the task-specific SFT. These findings reinforce the importance of supervised adaptation for certain downstream tasks. (4) Multi-source retrieval-augmented generation (RAG) offers competitive advantages. For instance, MRAG consistently achieves the highest recall, precision, and F1 score in convicting accuracy. However, while MRAG achieves high recall in referencing accuracy, its precision is relatively lower. We attribute this to our top-$k$ retrieval strategy (here $k=10$), which occasionally supplies multiple irrelevant statutes. As indicated by Table 2, only 4.31 criminal statutes are truly relevant per case. Since current LLMs are not fully adept at filtering out misleading articles, this results in a precision drop.